You have $21,600 to invest in a stock portfolio. Your choices are Stock X with an expected return of 14.3 percent and Stock Y with an expected return of 8.1 percent. Your goal is to create a portfolio with an expected return of 12.5 percent. All money must be invested. How much will you invest in Stock X?

A) $15,800

B) $18,273

C) $14,600

D) $15,329

E) $19,208
Answer #1

Ans D) $15,329

Let investment in Stock X be "x"
RETURN FROM X + RETURN FROM Y = REQUIRED RETURN
x * 14.3% + (21600 - x) * 8.1% = 21600 * 12.5%
0.143x + 1749.60 - 0.081x = 2700
0.062 x = 950.4
x = 15329.03